A father is 3 times as old as his son. 10 years later , father is 2 times as old as his son, What are their ages?
Answer by mananth(16946)

son's age x
father's age 3x
10 years later
son's age x+10
father's age 3x+10
father's age 2times older
3x+10=2(x+10)
3x+10=2 x+20
x=20-10
x=10 son's age
x+20=30 years father's age